OpenStudy (anonymous):

Consider the functions f(x) = 4x2 and g(x) = 5(x + 7). Find f[g(−5)]. Please show all of your work for full credit.

OpenStudy (amorfide):

g(x) = 5(x + 7) g(-5)=5(-5+7)=10 f(g)=f(10) f(10)=4(10)²=400

OpenStudy (anonymous):

thank you again @amorfide